# Leveraging Structural Knowledge in Diffusion Models for Source Localization in Data-Limited Graph Scenarios

### v2

@@ Abstract.
The source localization problem in graph information propagation is crucial for managing various network disruptions, from misinformation spread to infrastructure failures. While recent deep generative approaches have shown promise in this domain, their effectiveness is limited by the scarcity of real-world propagation data. This paper introduces SIDSL (Structure-prior Informed Diffusion model for Source Localization), a novel framework that addresses three key challenges in limited-data scenarios: unknown propagation patterns, complex topology-propagation relationships, and class imbalance between source and non-source nodes. SIDSL incorporates topology-aware priors through graph label propagation and employs a propagation-enhanced conditional denoiser with a GNN-parameterized label propagation module (GNN-LP). Additionally, we propose a structure-prior biased denoising scheme that initializes from structure-based source estimations rather than random noise, effectively countering class imbalance issues. Experimental results across four real-world datasets demonstrate SIDSL’s superior performance, achieving 7.5-13.3% improvements in F1 scores compared to state-of-the-art methods. Notably, when pretrained with simulation data of synthetic patterns, SIDSL maintains robust performance with only 10% of training data, surpassing baselines by more than 18.8%. These results highlight SIDSL’s effectiveness in real-world applications where labeled data is scarce.
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 2025; arxiv
@@ 1. Introduction
In today’s highly interconnected world, graph information propagation issues, such as misinformation spread, cyber threats, and infrastructure failures, have far-reaching consequences for society (Ding et al., 2024). The ability to quickly identify the source of these disruptions is critical for mitigating their impact. By analyzing snapshots of affected networks, we can trace the origin of the spread, a process essential for managing crises like disease outbreaks (Ru et al., 2023), enhancing network security (Kephart and White, 1993), and preventing further damage in scenarios such as power grid failures (Amin and Schewe, 2007).
@@
Early methods (Lappas et al., 2010; Shah and Zaman, 2012; Prakash et al., 2012; Luo et al., 2013; Zhu and Ying, 2014a, b) for source localization in graphs rely on metrics or heuristics derived from the graph’s topology, applicable only to specific propagation patterns like the Susceptible-Infected (SI) or Independent Cascade (IC) models. Notably, Wang et al. (Wang et al., 2017) overcome this limitation by introducing a label propagation algorithm based on the intuition of source centrality. Nevertheless, the approach has limitations in both expressiveness and scalability, as they cannot effectively encode graph topological information. Data-driven methods (Dong et al., 2019; Wang et al., 2022; Hou et al., 2023) overcome these limitations as they directly learn graph neural networks (GNNs) to capture the propagation process exhibited in empirical data, but still neglect the indeterminacy of information propagation that corresponds to the uncertain nature of source localization (Ling et al., 2022). Recently, deep generative models including variational autoencoders (Ling et al., 2022), normalization flows (Xu et al., 2024) and diffusion models (Huang et al., 2023b; Yan et al., 2024) have been adopted for solving the source localization problem. These probabilistic generative methods can quantify the indeterminacy in source localization by learning the empirical data distribution and advance the state-of-the-art performance.
However, collecting real-world propagation data for deep generative methods is difficult and costly, posing significant requirements on source localization models that can adapt to real-world environments with limited data. This brings up three main following challenges.
  • Firstly, real-world graphs typically exhibit unknown propagation patterns, which becomes far more challenging to characterize when data is limited. In this regard, existing methods (Dong et al., 2019; Wang et al., 2022; Ling et al., 2022; Yan et al., 2024) rely purely on data to gain an understanding of the propagation patterns, limiting their capability to generalize in unseen scenarios.
  • Secondly, complex interrelations between propagation patterns and graph topology are difficult to capture with limited data. Existing deep learning methods rely on a large amount of labeled data from the target network (i.e., identified source nodes from historical propagation) to account for the impact of structural heterogeneity on propagation patterns. Due to this heavy dependence on labeled data, these models typically underperform when applied to networks with limited training samples, hindering their practical applicability.
  • Thirdly, the inherent class imbalance between source and non-source nodes becomes more harmful under data scarcity, compromising the accuracy and reliability of source identification. In most propagation scenarios, source nodes naturally constitute a small minority of the total graph nodes (Cheng et al., 2024). Limited training data can further amplify this problem, leading to biased distribution models that tend to degenerate towards predicting all nodes as non-sources.
@@
Therefore, in this paper, we propose a novel generative diffusion framework, namely Structure-prior Informed Diffusion Model for Source Localization (SIDSL), which leverages topology-aware priors to achieve robust source localization with limited propagation data collected from real environments. Taking both graph topology and observed node states as input, SIDSL employs a denoising diffusion model guided by structural priors to predict source node distributions. To tackle the first challenge, We leverage graph structure-based source estimations, generated through graph label propagation, which identifies potential sources based on locally highest infection values. Integrated into the denoising network, these topology-aware priors provide stable guidance across different propagation patterns, enhancing generalization to unseen propagation scenarios. To address the second challenge, We design a propagation-enhanced conditional denoiser with a GNN-parameterized label propagation module (GNN-LP) that combines label propagation for tracing infection pathways with GNN for efficient topology-aware feature extraction, enabling effective learning of topology-propagation relationships from limited data. To tackle the third challenge, we develop a structure-prior biased denoising process that initializes denoising process from structure-based source estimations rather than random noise, creating a natural bias towards potential source nodes to prevent degeneration with limited, imbalanced data.
@@
The above designs work together in our structure-prior informed diffusion framework to capture stable structural patterns independent of specific propagation dynamics, which allows our model to learn pattern-invariant features using synthetic propagation data simulated from established models (Keeling and Eames, 2005; Kempe et al., 2003; Kermack and McKendrick, 1927). By focusing on these invariant features rather than specific propagation patterns, our model can effectively transfer the learned knowledge to real-world scenarios through efficient few-shot or zero-shot learning, as shown in Figure 1.
@@
(1) We propose a structure-prior informed diffusion framework that effectively addresses the challenge of limited data in source localization by incorporating topology-aware priors, enabling robust generalization to unknown propagation patterns in real-world scenarios.
(2) We introduce a series of innovative techniques, including a propagation-enhanced conditional denoiser with GNN-LP module and estimation-biased denoising, which work synergistically to handle structural heterogeneity and class imbalance issues.
(3) We evaluate SIDSL’s performance across four real-world datasets and demonstrate its superior effectiveness in source identification tasks. Our method consistently outperforms state-of-the-art baselines by 7.5-13.3% in F1 scores. Through effective pretraining with simulation data of synthetic patterns, SIDSL maintains robust performance with limited empirical data, surpassing baselines by over 19% in few-shot (using only 10% training samples) and 40% in zero-shot settings, demonstrating strong generalization capability in real-world applications. Additional evaluations further validate SIDSL’s effective synthetic-to-real transfer capability in improving sample efficiency and reducing training time.
Figure 1. The proposed SIDSL approach.
@@ 2.1. Source Localization
As the inverse problem of information propagation on networks, source localization refers to inferring the initial propagation sources given the current diffused observation, such as the states of the specified sensors or a snapshot of the whole network status (Shelke and Attar, 2019). It can be applied to tasks like rumor source identification and finding the origin of rolling blackouts in intelligent power grids (Shelke and Attar, 2019). Early approaches are rule-based and rely on metrics or heuristics derived from the network’s topology for source identification (Shah and Zaman, 2011; Zhu and Ying, 2014a, b). For example, Shah and Zaman (2011) develop a rumor-centrality-based maximum likelihood estimator under the Susceptible-Infected (SI) (Kermack and McKendrick, 1927) propagation pattern. This kind of method fails to effectively encode topology information. Later, deep learning-based methods devised for capturing the topological effect exhibited in empirical data have been proposed (Lappas et al., 2010; Luo et al., 2013; Wang et al., 2017; Dong et al., 2019; Wang et al., 2022; Hou et al., 2024). However, most of them fail to model the uncertainty of the location of sources, as the forward propagation process is stochastic. To overcome this, deep generative models have been adopted (Ling et al., 2022; Yan et al., 2024; Wang et al., 2024; Xu et al., 2024; Huang et al., 2023b). SLVAE (Ling et al., 2022) utilizes the Variational Auto-Encoders (VAEs) backbone and optimizes the posterior for better prediction. However, it is difficult to converge when the propagation pattern is complicated due to the nature of VAEs. DDMSL (Yan et al., 2024) models the Susceptible-Infected-Recovered (SIR)  (Kermack and McKendrick, 1927)infection process into the discrete Diffusion Model (DM) (Ho et al., 2020), and design a reversible residual block based on Graph Convolutional Networks (GCNs) (Kipf and Welling, 2016). However, it requires additional intermediate propagation data and cannot be generalized to other propagation patterns. Our method demonstrates superior functionality and adaptability for real-world applications, requiring fewer input data while addressing existing limitations, thus offering greater practical value. We provide a comparison of typical source localization methods in the Appendix A.
@@ 2.2. Typical Propagation Models
Information propagation estimation models information spread in networks and explains propagation sources, with applications in event prediction (Zhao, 2021), adverse event detection (Wang and Zhao, 2018), and disease spread prediction (Tang et al., 2023). Two main model categories exist: infection models and influence models. Infection models like Susceptible-Infected (SI) and Susceptible-Infected-Susceptible (SIS) manage transitions between susceptible and infected states in networks (Kermack and McKendrick, 1927; Keeling and Eames, 2005). In these models, infected nodes attempt to infect adjacent nodes with probability $\beta$ at each iteration, while in SIS, infected nodes may revert to susceptible with probability $\lambda$. The Susceptible-Infected-Recovered (SIR) model extends this by adding a recovered state.
Independent Cascade (IC) and Linear Threshold (LT) (Kempe et al., 2003) are influence models that examine influence spread in social and infrastructure networks. In the IC model, nodes are either active or inactive, starting with initial active nodes. Newly activated nodes get one chance to activate inactive neighbors, with activation probability based on edge weight. The LT model activates inactive nodes when accumulated neighbor influence exceeds a threshold.
@@ 3.1. Problem Formulation
Our research problem is formulated as follows. Given an undirected social network $\mathcal{G}=(V,E)$ where $V$ is the node set, $E$ is the edge set, and $Y=\{Y_{1},\ldots,Y_{|V|}\}$ is an infection state of all nodes in $\mathcal{G}$, which describes that a subset of nodes in $\mathcal{G}$ have been infected. Each $Y_{i}\in\{1,0\}$ denotes the infection state of node $v_{i}\in V$, where $Y_{i}=1$ indicates that $v_{i}$ is infected and otherwise $Y_{i}=0$ indicates it is uninfected. We aim to find the original propagation source $\hat{X}$ from the propagated observation $Y$, so that the loss with the ground Truth source set $X^{*}\in\{1,0\}^{|V|\times 1}$ is minimized, i.e. $\hat{X}=argmin_{X}||X-X^{*}||_{2}^{2}$. To account for the uncertainty in source localization, we need to construct a probabilistic model $P(X|Y,\mathcal{G})$, which can be used to sample for the final prediction.
@@ 3.2. Label Propagation based Source Identification
In realistic situations, the intractable propagation process does not have an explicit prior, and it is also challenging to value appropriate parameters for the pre-selected underlying propagation model. To address this, Wang et al. (2017) propose Label Propagation based Source Identification (LPSI). Since LPSI investigates the same problem as our research, we use it as a baseline to compare performance. LPSI captures source centrality characteristics in the method design. The centrality of sources shows that nodes far from the source are less likely to be infected than those near it (Shah and Zaman, 2012), which can also be observed in the real-world data by our analysis in the Appendix B. Based on these ideas, they propose to perform label propagations on the observation state of the network. By setting $Y[Y=0]=-1$ and ${}\mathcal{Z}^{t=0}\xleftarrow{}Y$, the iteration of label propagation and the convergence states are as follows:
@@
$\mathcal{Z}$ finally converges to:$\mathcal{Z}^{*}=(1-\alpha)(I-\alpha S)^{-1}Y$,where $S=D^{-1/2}AD^{-1/2}$ is the normalized weight matrix of graph $\mathcal{G}$, $\alpha$ is the fraction of label information from neighbors, and $\mathcal{N}(i)$ stands for the neighbor set of the node $i$. After obtaining the converged label matrix $\mathcal{Z}^{*}$, one node can identified as a source when its final label is larger than its neighbors. While node labels alone cannot fully capture complex structural information, this method still effectively identifies structural patterns related to source centrality (Appendix B). In our work, we leverage the idea of LPSI to generate structural guidance for efficient source identification and details can be found in Section 4.
@@ 4.1. Diffusion Model for Source Localization
To capture the indeterminacy of the ill-posed localization problem, it is essential to build a probabilistic model that can also leverage the topological information in the graph structure. We consider using the generative diffusion model (DM) framework to tackle this requirement by modifying it as a source predictor, which classifies each node into two categories: source or non-source. The diffusion model aims to learn source distributions by gradually adding Gaussian noise in samples in the forward process and learning to reverse this process using denoising networks conditioned on observations and graph structure.
@@
where $\beta_{t}$ is the noise schedule, $X_{0}$ represents the initial source labels, $Y$ denotes the observations, $\mathcal{G}$ represents the graph structure, and $X_{t}$ denotes the noisy features at timestep $t$. At the end of forward process, $X_{n}$ becomes pure Gaussian noise. While the forward process only requires source labels, the reverse process leverages observation $Y$ and graph structure $\mathcal{G}$ as conditional inputs to guide the denoising.
In the reverse process, it gradually transforms pure Gaussian noise into the original source labels, generating new samples $\hat{X}\in[0,1]^{|V|\times 1}$ representing the probability of each node being a source. The final source predictions are obtained by thresholding these probabilities. To enhance stability across different propagation patterns, we leverage graph structure-based source estimations as conditional priors. These priors $X_{est}=\mathcal{Z}^{*}$ are obtained through label propagation (Section 3.2), which identifies potential sources by aggregating infection values through the graph’s normalized weight matrix while capturing source centrality characteristics. The convergence result $\mathcal{Z}^{*}$ reflects each node’s structural importance relative to the observed infection pattern.
@@
where $\mu_{\theta}$ is parameterized by a denoising network $f_{\theta}$ that predicts the mean of the Gaussian distribution and $\sigma_{t}$ is the predicted variance. The diffusion framework, whose iterative diffusion process enables fine-grained integration of prior knowledge at multiple scales through its progressive denoising steps, allows for elegant integration of prior knowledge. By integrating topology-aware priors $X_{est}$ as stable guidance across different propagation patterns, our diffusion framework combines structural knowledge with data-driven pattern learning, enhancing generalization to new propagation scenarios.

### v3

# Structure-prior Informed Diffusion Model for Graph Source Localization with Limited Data

@@ Abstract.
Source localization in graph information propagation is essential for mitigating network disruptions, including misinformation spread, cyber threats, and infrastructure failures. Existing deep generative approaches face significant challenges in real-world applications due to limited propagation data availability. We present SIDSL (Structure-prior Informed Diffusion model for Source Localization), a generative diffusion framework that leverages topology-aware priors to enable robust source localization with limited data. SIDSL addresses three key challenges: unknown propagation patterns through structure-based source estimations via graph label propagation, complex topology-propagation relationships via a propagation-enhanced conditional denoiser with GNN-parameterized label propagation module, and class imbalance through structure-prior biased diffusion initialization. By learning pattern-invariant features from synthetic data generated by established propagation models, SIDSL enables effective knowledge transfer to real-world scenarios. Experimental evaluation on four real-world datasets demonstrates superior performance with 7.5-13.3% F1 score improvements over baselines, including over 19% improvement in few-shot and 40% in zero-shot settings, validating the framework’s effectiveness for practical source localization. Our code can be found here.
Diffusion Model; Graph Source Localization; Information Diffusion
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761146isbn: 979-8-4007-2040-6/2025/11ccs: Networks Network dynamicsccs: Information systems Social networksccs: Information systems Data mining
@@ 1. Introduction
Graph information propagation issues, such as misinformation spread, cyber threats, and infrastructure failures, have far-reaching societal consequences. Quickly identifying disruption sources is critical for impact mitigation. By analyzing snapshots of affected networks, we can trace spread origins, essential for managing disease outbreaks (Ru et al., 2023; Ding et al., 2024), network security (Kephart and White, 1993), and infrastructure failures (Amin and Schewe, 2007).
@@
Early methods (Lappas et al., 2010; Shah and Zaman, 2012; Prakash et al., 2012; Luo et al., 2013; Zhu and Ying, 2014a, b) for source localization in graphs rely on topology-derived metrics applicable only to specific propagation patterns like the Susceptible-Infected (SI) or Independent Cascade (IC) models. Wang et al. (Wang et al., 2017) introduced label propagation based on source centrality, but with limited expressiveness and scalability in encoding topological information. Data-driven methods (Dong et al., 2019; Wang et al., 2022; Hou et al., 2023) overcome these limitations as they directly learn graph neural networks (GNNs) to capture propagation processes but neglect propagation indeterminacy (Ling et al., 2022). Recent deep generative models, including variational autoencoders (Ling et al., 2022), normalization flows (Xu et al., 2024), and diffusion models (Huang et al., 2023b; Yan et al., 2024), quantify this indeterminacy by learning empirical data distributions, advancing state-of-the-art performance.
However, collecting real-world propagation data for deep generative methods is difficult and costly, requiring models that adapt to limited data environments. This poses three main challenges. First, real-world graphs exhibit unknown propagation patterns that are challenging to characterize with limited data. Existing methods (Dong et al., 2019; Wang et al., 2022; Ling et al., 2022; Yan et al., 2024) rely purely on data for pattern understanding, limiting generalization to unseen scenarios. Second, complex interrelations between propagation patterns and graph topology are difficult to capture with limited data. Existing methods require extensive labeled data from target networks to account for the structural heterogeneity’s impact on propagation patterns. This dependence causes underperformance with limited training samples, hindering practical applicability. Third, inherent class imbalance between source and non-source nodes becomes more harmful under data scarcity, compromising identification accuracy. This challenge remains unaddressed by existing solutions and requires comprehensive handling of both abundant and limited data scenarios through novel approaches that prevent model degeneration towards non-source predictions.
Figure 1. The proposed SIDSL approach.
Therefore, in this paper, we propose a novel generative diffusion framework, namely Structure-prior Informed Diffusion Model for Source Localization (SIDSL), which leverages topology-aware priors to achieve robust source localization with limited propagation data collected from real environments. SIDSL employs a denoising diffusion model guided by structural priors to predict source node distributions from graph topology and observed node states.
@@
To tackle the first challenge, we leverage graph structure-based source estimations, generated through graph label propagation, which identify potential sources based on locally highest infection values. Integrated into the denoising network, these topology-aware priors provide stable guidance across different propagation patterns, enhancing generalization to unseen propagation scenarios. To address the second challenge, we design a propagation-enhanced conditional denoiser with a GNN-parameterized label propagation module (GNN-LP) that combines label propagation for tracing infection pathways with GNN for efficient topology-aware feature extraction, enabling effective learning of topology-propagation relationships from limited data. To tackle the third challenge, we develop a structure-prior biased diffusion process that comprehensively handles the fundamental source/non-source imbalance by initializing from structure-based estimations, providing a complete solution that advances imbalance resolution in both data-rich and data-scarce scenarios.
@@
The above designs work together in our structure-prior informed diffusion framework to capture stable structural patterns independent of specific propagation dynamics, allowing our model to learn pattern-invariant features using synthetic propagation data simulated from established models (Keeling and Eames, 2005; Kempe et al., 2003; Kermack and McKendrick, 1927; Pan et al., 2025). This allows effective knowledge transfer to real-world scenarios through few-shot or zero-shot learning (Figure 1).
@@
(1) We propose SIDSL, a structure-prior informed diffusion framework that uniquely integrates structural priors with diffusion models to effectively address the challenge of limited labeled data in source localization. Unlike existing approaches requiring abundant training data, our framework leverages domain knowledge and topology-aware priors for robust generalization to unknown propagation patterns in real-world low-data scenarios.
(2) We introduce a series of innovative techniques, including a propagation-enhanced conditional denoiser with GNN-LP module and structure-prior biased denoising, which work synergistically to handle structural heterogeneity and class imbalance issues.
(3) We evaluate SIDSL’s performance across four real-world datasets and demonstrate its superior effectiveness in source identification tasks. Our method achieves superior performance across four real-world datasets, outperforming baselines by 7.5-13.3% in F1 scores, with over 19% improvement in few-shot and 40% in zero-shot settings. Additional evaluations further validate SIDSL’s effective synthetic-to-real transfer capability in improving sample efficiency and reducing training time.
@@ 2.1. Source Localization
Source localization, the inverse problem of information propagation in networks, infers initial sources from diffused observations, with applications such as rumor source identification and power grid failure detection (Shelke and Attar, 2019; Su et al., 2024). Early approaches are rule-based and rely on metrics or heuristics derived from the network’s topology for source identification (Shah and Zaman, 2011; Zhu and Ying, 2014a, b). For instance, Shah and Zaman (2011) developed a rumor-centrality-based maximum likelihood estimator for the Susceptible-Infected (SI) (Kermack and McKendrick, 1927) model. These methods, however, inadequately encode topology. Subsequently, deep learning methods emerged (Lappas et al., 2010; Luo et al., 2013; Wang et al., 2017; Dong et al., 2019; Wang et al., 2022; Hou et al., 2024), but most struggled to model source location uncertainty in stochastic propagation. To overcome this, deep generative models have been adopted (Ling et al., 2022; Yan et al., 2024; Wang et al., 2024; Xu et al., 2024; Huang et al., 2023b). SLVAE (Ling et al., 2022) utilizes Variational Auto-Encoders (VAEs) and optimizes the posterior for better prediction. However, VAEs struggle with complex propagation patterns. DDMSL (Yan et al., 2024) models the Susceptible-Infected-Recovered (SIR)  (Kermack and McKendrick, 1927)infection process into Diffusion Models (DM) (Ho et al., 2020; Xu et al., 2025), and designs a reversible residual block based on Graph Convolutional Networks (GCNs) (Kipf and Welling, 2016). However, it requires additional intermediate propagation data and cannot be generalized to other propagation patterns. In contrast to existing methods that demand extensive training data, our novel framework leverages domain knowledge for robust localization with minimal labeled data—a significant advance for real-world scenarios. A comparison of typical methods is in Appendix B.
@@ 2.2. Typical Propagation Models
Information propagation models estimate information spread in networks, applied in event prediction (Zhao, 2021), adverse event detection (Wang and Zhao, 2018), disease spread prediction (Tang et al., 2023), and so on. Two main categories exist: infection models and influence models. Infection models such as Susceptible-Infected (SI) and Susceptible-Infected-Susceptible (SIS) manage state transitions in networks (Kermack and McKendrick, 1927; Keeling and Eames, 2005). Infected nodes attempt to infect neighbors with probability $\beta$, while in SIS models, infected nodes revert to susceptible with probability $\lambda$. The Susceptible-Infected-Recovered (SIR) model adds a recovered state. Influence models include Independent Cascade (IC) and Linear Threshold (LT) (Kempe et al., 2003). In IC, newly activated nodes attempt once to activate neighbors based on edge weights. LT models activate nodes when accumulated neighbor influence exceeds a threshold.
@@ 3.1. Problem Formulation
[…]
@@ 3.2. Label Propagation based Source Identification
In realistic situations, the intractable propagation process does not have an explicit prior, and it is also challenging to value appropriate parameters for the pre-selected underlying propagation model. To address this, Wang et al. (2017) propose Label Propagation based Source Identification (LPSI). Since LPSI investigates the same problem as our research, we use it as a baseline to compare performance. LPSI captures source centrality characteristics in the method design. The centrality of sources shows that nodes far from the source are less likely to be infected than those near it (Shah and Zaman, 2012), which can also be observed in the real-world data by our analysis in the Appendix E. Based on these ideas, they propose to perform label propagation on the observation state of the network. By setting Y[Y=0]=1Y[Y=0]=-1 and $\mathcal{Z}^{t=0}\xleftarrow{}Y$, the iteration of label propagation and the convergence states are as follows:
@@
[…]
@@ 4.1. Diffusion Model for Source Localization
To capture the indeterminacy of the ill-posed localization problem, it is essential to build a probabilistic model that can also leverage the topological information in the graph structure. We consider using the generative diffusion model (DM) framework to tackle this requirement by modifying it as a source predictor, which classifies nodes as source or non-source. The diffusion model aims to learn source distributions by gradually adding Gaussian noise in samples in the forward process and learning to reverse this process using denoising networks conditioned on observations and graph structure.
@@
where $\beta_{t}$ is the noise schedule, $X_{0}$ the initial source labels, $Y$ the observations, $\mathcal{G}$ the graph structure, and $X_{t}$ the noisy features at timestep $t$. At $t=n$, $X_{t}$ becomes pure Gaussian noise, while the reverse process leverages $Y$ and $\mathcal{G}$ to guide denoising.
In the reverse process, it gradually transforms pure Gaussian noise into the original source labels, generating samples $\hat{X}\in[0,1]^{|V|\times 1}$ representing source probabilities, with final predictions obtained by thresholding. To enhance stability across different propagation patterns, we leverage graph structure-based source estimations as conditional priors. These priors $X_{est}=\mathcal{Z}^{*}$, standing for potential sources, are obtained through label propagation (Section 3.2), aggregating infection values through the graph’s normalized weight matrix to capture source centrality and structural importance.
@@
where $\mu_{\theta}$ is parameterized by a denoising network $f_{\theta}$ that predicts the mean of the Gaussian distribution and $\sigma_{t}$ is the predicted variance. The iterative diffusion process enables fine-grained prior knowledge integration through progressive denoising. By incorporating topology-aware priors $X_{est}$, our framework combines structural knowledge with data-driven learning, enhancing generalization across propagation scenarios.